# The Nondecreasing Rank

### v1

@@ 10 Finding Low ND Rank Approximations
If $\mathbf{T}$} has an optimal, unconstrained, rank-$r$ decomposition ${\mathbf{T}}^{(r)}$ with ${\mathbf{T}}^{(r)}\in\mathcal{N}_{\leq r}$ then ${\mathbf{T}}^{(r)}=\mathrm{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq r}}\;\|{\mathbf{T}}-\bm{\theta}\|_{F}^{2}$. For matrices with ${\mathbf{T}}^{(2)}\in\mathcal{N}_{<\infty}$ we have that ${\mathbf{T}}^{(2)}=\mathrm{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq 2}}\;\|{\mathbf{T}}-\bm{\theta}\|_{F}^{2}$.
@@
There are a wide variety of algorithms for computing approximate nonnegative factorizations, the two main classes of which are multiplicative update (MU) algorithms and least squares algorithms; see [28, Ch 8] for a comprehensive account. Below we propose an algorithm [15] for minimizing the Frobenius norm between a data tensor and a low ND rank tensor. This procedure is related to the hierarchical alternating least squares (HALS) algorithm and is based off minimizing the following expression
@@
with respect to ${\mathbf{v}}^{(st)}\in\mathcal{C}(\mathcal{P}_{t})$, while holding all other vectors fixed. If
@@
is the unconstrained minimizer of $v_{l}^{(st)}$, the objective function in (11) can be written as
$\sum_{l=1}^{p_{t}}\bigg{\|}\tilde{{\mathbf{T}}}_{\bullet\cdots\bullet l\bullet\cdots\bullet}-\otimes_{j\neq t}{\mathbf{v}}^{(sj)}\tilde{v}_{l}^{(st)}\bigg{\|}_{F}^{2}+\bigg{(}\prod_{j\neq t}\|{\mathbf{v}}^{(sj)}\|_{2}^{2}\bigg{)}\sum_{l=1}^{p_{t}}(\tilde{v}_{l}^{(st)}-v_{l}^{(st)})^{2},$
where only the second term involves ${\mathbf{v}}^{(st)}$. Importantly, this second term is separable in $v_{l}^{(st)}$, $l\in[p_{t}]$, indicating that the efficient, pool-adjacent-violators algorithm (PAVA) [18, 54] can be used to minimize this quadratic program over $\mathcal{C}(\mathcal{P}_{t})$. Our ND hierarchical least squares algorithm is a block-coordinate descent routine that proceeds by cycling through the PAVA updates for each vector ${\mathbf{v}}^{(ij)}$ until convergence.
@@
Due to the Frobenius norm optimization problem being jointly non-convex in the ${\mathbf{v}}^{(ij)}$s there is no guarantee that the sequence of updates converges to a global optimum. However, Proposition 3.7.1 of [9] ensures that any limit point of the ${\mathbf{v}}^{(ij)}$ iterates is a stationary point. As there can exist multiple stationary points, it is recommended that multiple, distinct initializations be chosen in the ND HALS algorithm. One possible choice of initialization is to first run an unconstrained, rank-$r$ approximation algorithm on $\mathbf{T}$}, such as alternating least squares [35], to obtain vectors $\tilde{{\mathbf{v}}}^{(ij)}$. These vectors can then respectively be orthogonally projected onto the cone $\mathcal{C}(\mathcal{P}_{j})$ by again using PAVA to solve ${\mathbf{v}}^{(ij)}=\text{argmin}_{{\mathbf{v}}^{(ij)}\in\mathcal{C}(\mathcal{P}_{j})}\|\tilde{{\mathbf{v}}}^{(ij)}-{\mathbf{v}}^{(ij)}\|_{2}^{2}$.
@@ 11.1 Pig Weight Data
The pig weight dataset [19, Sec 3.2] available in the fds R package [47] records the weights of $48$ growing pigs over the course of nine weeks. Figure 5(a) illustrates that all of the pigs have monotonically increasing weights over time. If ${\mathbf{T}}\in\mathbb{R}^{48\times 9}$ is the matrix of weights we may assume that the poset corresponding to the columns is a chain, while no ordering constraints, apart from nonnegativity, are placed on the rows. If ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}^{(i)}({\mathbf{b}}^{(i)})^{\intercal}$ is a rank-$r$ ND factorization, the weight profile of pig $j$ is ${\mathbf{T}}_{j\cdot}=\sum_{i=1}^{r}a^{(i)}_{j}{\mathbf{b}}^{(i)}$. Up to scale factors, the vectors ${\mathbf{b}}^{(i)}$ can be interpreted as growth curves for $r$ different “subpopulations” of pigs, while $a^{(i)}_{j}$ indicates the degree of membership of pig $j$ to subpopulation $i$. The growth curve of each pig is modeled to approximately be a conical combination of the subpopulation curves. If only an NMF was applied to this dataset the ${\mathbf{b}}^{(i)}$ may not be as easily interpretable as growth curves. Furthermore, the second, right, singular vector certainly cannot be interpreted as a growth curve as it does not even contain entries that have the same sign.
@@
To fit this dataset we find an ND rank two approximation ${\mathbf{T}}\approx{\mathbf{a}}^{(1)}\otimes{\mathbf{b}}^{(1)}+{\mathbf{a}}^{(2)}\otimes{\mathbf{b}}^{(2)}$ with ${\mathbf{a}}^{(i)}\in\mathbb{R}^{48}_{+}$ and ${\mathbf{b}}^{(i)}\in\mathcal{C}(\{1,\ldots,9\})$. A rank of two was chosen because this rank is sufficient to yield a low-reconstruction error for $\mathbf{T}$}; the first four squared singular value of $\mathbf{T}$} are 121,000121{,}000, $613$, $129$, and $109$. The rank-two SVD approximation ${\mathbf{T}}^{(2)}$ of $\mathbf{T}$} has nonnegative entries and monotone increasing rows, implying that ${\mathbf{T}}^{(2)}\in\mathcal{N}_{\leq 2}$ (Theorem 12). The vectors ${\mathbf{b}}_{1},{\mathbf{b}}_{2}$ can be taken to be the pair of rows of ${\mathbf{T}}^{(2)}$ that have the largest angle between each other. Such a pair can easily be found by computing dots products between the rows of ${\mathbf{T}}^{(2)}$ that are rescaled to have unit norm. The corresponding coefficients ${\mathbf{a}}_{1},{\mathbf{a}}_{2}$ are found by a linear regression. Lastly, we rescale ${\mathbf{b}}_{1},{\mathbf{b}}_{2}$ so that the mean weight of each ${\mathbf{b}}^{(i)}$ is equal to the average weight across all the pigs: $\tfrac{1}{9}\sum_{l=1}^{9}b^{(i)}_{l}=\tfrac{1}{9*48}\sum_{i=1}^{48}\sum_{j=1}^{9}T_{ij}$. Figure 5(a) displays the two estimated growth curves ${\mathbf{b}}^{(1)}$ and ${\mathbf{b}}^{(2)}$, overlaid on the observed growth curves. Interestingly, around week five the slope of both estimated curves change, a phenomena that may warrant further study. In Figure 5(b) we see that the fitted curve $a_{i}^{(1)}{\mathbf{b}}^{(1)}+a_{i}^{(2)}{\mathbf{b}}^{(2)}$, for pig $i$, is nearly a convex combination of ${\mathbf{b}}^{(1)}$ and ${\mathbf{b}}^{(2)}$, as the scale factor $a^{(1)}_{i}+a^{(2)}_{i}$ is close to one. The relative importance of ${\mathbf{b}}^{(1)}$ as compared to ${\mathbf{b}}^{(2)}$ does noticeably vary across the pigs.
@@
One issue that has not yet been discussed in this work is that ND factorizations are not unique, even after accounting for scaling and permutation of the rank-one factors. Various regularization penalties or additional constraints have been introduced in the NMF literature [28, Ch 4] to select a factorization that has desirable properties. In the above analysis we have implicitly selected the minimum volume [39] ND factorization by choosing rows from ${\mathbf{T}}^{(2)}$ to serve as the ${\mathbf{b}}^{(i)}$. An alternative here is take a maximum volume factorization by choosing the ${\mathbf{b}}^{(i)}$ to be the extremal rays of the cone $\mathcal{C}(\{1,\ldots,9\})\cap\text{row}({\mathbf{T}}^{(2)})$. Either factorization will yield the same reconstruction error $\|{\mathbf{T}}-{\mathbf{T}}^{(2)}\|^{2}_{F}$, but will lead to potentially different interpretations with respect to the ${\mathbf{b}}^{(i)}$ curves and the ${\mathbf{a}}^{(i)}$ membership weights.
@@ 11.2 Perceived Mental Health of Canadians
We examine a subset of the data from the annual Canadian Community Health Survey (CCHS) [51] from 20162016-$2022$ that records responses of whether an individual perceives their mental health to be either fair or poor. The survey also records the gender and age group of the participant. The $(5,7,2)$-dimensional tensor $\mathbf{T}$} in Table 2 displays the proportion of respondents who indicated that they had either poor or fair mental health.
@@
Using the ND HALs algorithm from Section 10 we fit both rank-two and rank-one approximations to the data tensor of the form $\sum_{i=1}^{r}{\mathbf{a}}^{(i)}\otimes{\mathbf{b}}^{(i)}\otimes{\mathbf{c}}^{(i)}$. The residual sum of squares for the respective approximations are $145$ and $55$ with the total sum of squares equaling $6925$. We only impose nonnegativity constraints on the gender mode, while the Hasse diagrams in Figure 6 provides orderings on the age group and years that are consistent with the observed data. Specifically, the mental health of respondents deteriorated in 2020 and the subsequent two years, conceivably as a result of the COVID-19 pandemic. Respondents in the second youngest age group were most likely to perceive their mental health as fair or poor while respondents in the oldest age group were most likely to not provide this response.
There is a marked difference between how individuals in the 12-17 age group responded between females and males, with females having poorer perceived mental health. In fitting a rank-two ND decomposition the two rank-one factors provide a low-dimensional summary of this gender-age interaction. The rank-one factors of the two rank-one terms are shown in Table 3. These factors are scaled so that $\|{\mathbf{c}}^{(i)}\|_{1}=1$ and $\|{\mathbf{b}}^{(i)}\|_{1}=7$. We see that the factor ${\mathbf{a}}^{(2)}\otimes{\mathbf{b}}^{(2)}\otimes{\mathbf{c}}^{(2)}$ only alters the female portion of the tensor and represents a deviation between the male and female responses. In comparison with ${\mathbf{a}}^{(1)}$, the 12-17 age group in ${\mathbf{a}}^{(2)}$ takes on larger values relative to the 18-34 age group. Furthermore, the impact of pandemic is more apparent in ${\mathbf{b}}^{(2)}$ as compared to ${\mathbf{b}}^{(1)}$. In summary, the low ND rank decomposition allows for a concise interpretation of the data tensor and illustrates that young females in 2022 were particularly prone to respond that they had fair or poor mental health as compared to young males.
@@
Gender ${\mathbf{a}}^{(1)}$ ${\mathbf{a}}^{(2)}$
@@ 12 Conclusion and Future Directions
In this article the concept of the nondecreasing rank was introduced and properties of nondecreasing factorizations were examined. There remains a wealth of relevant, open questions to be explored, a few of which we mention here. One property of ND factorizations not extensively discussed is the uniqueness of such factorizations. Like with nonnegative matrix factorizations, it is expected that only under similarly strong conditions, such as separability [28, Sec 4.2.2], will a matrix have a unique ND factorization. Also not discussed, is how to choose a suitable rank $r$ for performing an ND factorization. An easy method for doing this is to plot the value of a divergence objective function against $r$ and look for a kink in the plot where the divergence stops decreasing rapidly in response to an increase of $r$ [14]. In Section 10 an algorithm is introduced for minimizing the Frobenius norm between a tensor and its ND approximation. Developing efficient multiplicative update [34] algorithms for finding ND approximations would be a valuable contribution. To construct such an algorithm more constraints as compared to NMF, stemming from the monotonicity constraints of the order cones, have to be contended with.
Nondecreasing factorizations are closely related to ideas in the field of order constrained statistical inference [6, 46]. Beyond monotonicity constraints, related shape constraints for functions, such as convexity and unimodality, would be interesting to inspect for low-rank structures. Furthermore, such problems can be contextualized from an infinite-dimensional, functional data analysis perspective, where notions of convex sets and rank continue to be well-defined.
13 Acknowledgements

### v2

@@ 10 Finding Low ND Rank Approximations
If $\mathbf{T}$} has an optimal, unconstrained, rank-$r$ decomposition ${\mathbf{T}}^{(r)}=\text{argmin}_{\bm{\theta}:\mathrm{rank}(\bm{\theta})\leq r}\|{\mathbf{T}}-\bm{\theta}\|_{F}^{2}$ where ${\mathbf{T}}^{(r)}\in\mathcal{N}_{\leq r}$ then ${\mathbf{T}}^{(r)}=\mathrm{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq r}}\;\|{\mathbf{T}}-\bm{\theta}\|_{F}^{2}$. For matrices with ${\mathbf{T}}^{(2)}\in\mathcal{N}_{<\infty}$ we have that ${\mathbf{T}}^{(2)}=\mathrm{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq 2}}\;\|{\mathbf{T}}-\bm{\theta}\|_{F}^{2}$.
@@
There are a wide variety of algorithms for computing approximate nonnegative factorizations, the two main classes of which are multiplicative update (MU) algorithms and least squares algorithms; see [29, Ch 8] for a comprehensive account. Below we propose an algorithm for minimizing the Frobenius norm between a data tensor and a low ND rank tensor. This procedure is closely related to the hierarchical alternating least squares (HALS) algorithm [16] and is based off minimizing the following expression
@@
with respect to ${\mathbf{v}}^{(st)}\in\mathcal{C}(\mathcal{P}_{t})$, while holding all other vectors fixed. We have defined $\tilde{{\mathbf{T}}}\coloneq{\mathbf{T}}-\sum_{i\neq s}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}$ in (11). Defining
@@
to be the unconstrained minimizer of $v_{l}^{(st)}$, the objective function in (11) can be written as
$\sum_{l=1}^{p_{t}}\bigg\|\tilde{{\mathbf{T}}}_{\bullet\cdots\bullet l\bullet\cdots\bullet}-\otimes_{j\neq t}{\mathbf{v}}^{(sj)}\tilde{v}_{l}^{(st)}\bigg\|_{F}^{2}+\bigg(\prod_{j\neq t}\|{\mathbf{v}}^{(sj)}\|_{2}^{2}\bigg)\sum_{l=1}^{p_{t}}(\tilde{v}_{l}^{(st)}-v_{l}^{(st)})^{2},$ (12)
where only the second term involves ${\mathbf{v}}^{(st)}$. This decomposition arises from a sum of squares decomposition in linear regression: $\|{\mathbf{y}}-{\mathbf{x}}\beta\|^{2}_{F}=\|{\mathbf{y}}-{\mathbf{x}}\hat{\beta}\|^{2}_{F}+\|{\mathbf{x}}\hat{\beta}-{\mathbf{x}}\beta\|^{2}_{F}$, with the roles of $^{\mathbf{y}},{\mathbf{x}},\hat{\beta}$, and $\hat{\beta}$ respectively being played by $\tilde{{\mathbf{T}}}_{\bullet\cdots\bullet l\bullet\cdots\bullet},\otimes_{j\neq t}{\mathbf{v}}^{(sj)},\tilde{v}_{l}^{(st)}$, and $v_{l}^{(st)}$. Importantly, the second term in (12) is separable in $v_{l}^{(st)}$, $l\in[p_{t}]$, indicating that the efficient, pool-adjacent-violators algorithm (PAVA) [19, 57] can be used to minimize this quadratic program over $\mathcal{C}(\mathcal{P}_{t})$. Our ND hierarchical least squares algorithm is a block-coordinate descent routine that proceeds by cycling through the PAVA updates for each vector ${\mathbf{v}}^{(ij)}$ until convergence.
@@
Due to the Frobenius norm optimization problem being jointly non-convex in the ${\mathbf{v}}^{(ij)}$s there is no guarantee that the sequence of updates converges to a global optimum. However, Proposition 2.7.1 of [9] ensures that any limit point of the ${\mathbf{v}}^{(ij)}$ iterates is a stationary point as long as none of the vectors ${\mathbf{v}}^{(ij)}$ are updated to the zero vector over the course of the algorithm. The issue that can occur when any ${\mathbf{v}}^{(sj)}={\mathbf{0}}$ is that the componentwise objective function (12) is not strictly convex because the relevant quadratic term vanishes. As there can exist multiple stationary points, it is recommended that multiple, distinct initializations be chosen in the ND HALS algorithm. One possible choice of initialization is to first run an unconstrained, rank-$r$ approximation algorithm on $\mathbf{T}$}, such as alternating least squares [36], to obtain vectors $\tilde{{\mathbf{v}}}^{(ij)}$. These vectors can then respectively be orthogonally projected onto the cone $\mathcal{C}(\mathcal{P}_{j})$ by again using PAVA to solve ${\mathbf{v}}^{(ij)}=\text{argmin}_{{\mathbf{v}}^{(ij)}\in\mathcal{C}(\mathcal{P}_{j})}\|\tilde{{\mathbf{v}}}^{(ij)}-{\mathbf{v}}^{(ij)}\|_{2}^{2}$.
@@ 11.1 Pig Weight Data
The pig weight dataset [20, Sec 3.2] available in the fds R package [48] records the weights of $48$ growing pigs over the course of nine weeks. Figure 5(a) illustrates that all of the pigs have monotonically increasing weights over time. If ${\mathbf{T}}\in\mathbb{R}^{48\times 9}$ is the matrix of weights we may assume that the poset corresponding to the columns is a chain, while no ordering constraints, apart from nonnegativity, are placed on the rows. If ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}^{(i)}({\mathbf{b}}^{(i)})^{\intercal}$ is a rank-$r$ ND factorization, the weight profile of pig $j$ is ${\mathbf{T}}_{j\cdot}=\sum_{i=1}^{r}a^{(i)}_{j}({\mathbf{b}}^{(i)})^{\intercal}$. Up to scale factors, the vectors ${\mathbf{b}}^{(i)}$ can be interpreted as growth curves for $r$ different “subpopulations” of pigs, while $a^{(i)}_{j}$ indicates the degree of membership of pig $j$ to subpopulation $i$. The growth curve of each pig is modeled to approximately be a conical combination of the subpopulation curves. If only an NMF was applied to this dataset the ${\mathbf{b}}^{(i)}$ may not be as easily interpretable as growth curves. Furthermore, the second, right, singular vector certainly cannot be interpreted as a growth curve as it does not even contain entries that have the same sign.
@@
To fit this dataset we find an ND rank two approximation ${\mathbf{T}}\approx{\mathbf{a}}^{(1)}\otimes{\mathbf{b}}^{(1)}+{\mathbf{a}}^{(2)}\otimes{\mathbf{b}}^{(2)}$ with ${\mathbf{a}}^{(i)}\in\mathbb{R}^{48}_{+}$ and ${\mathbf{b}}^{(i)}\in\mathcal{C}(\{1,\ldots,9\})$. A rank of two was chosen because this rank is sufficient to yield a low-reconstruction error for $\mathbf{T}$}; the first four squared singular values of $\mathbf{T}$} are 121,000121{,}000, $613$, $129$, and $109$. The rank-two SVD approximation ${\mathbf{T}}^{(2)}$ of $\mathbf{T}$} has nonnegative entries and monotone increasing rows, implying that ${\mathbf{T}}^{(2)}\in\mathcal{N}_{\leq 2}$ (Theorem 11). The vectors ${\mathbf{b}}_{1},{\mathbf{b}}_{2}$ can be taken to be the pair of rows of ${\mathbf{T}}^{(2)}$ that have the largest angle between each other. Such a pair can easily be found by computing dot products between the rows of ${\mathbf{T}}^{(2)}$ that are rescaled to have unit norm. This approach has been used in the NMF literature for matrices that have a nonnegative rank of two which have the property of being separable [29, Sec 4.1]. The corresponding coefficients ${\mathbf{a}}^{(1)},{\mathbf{a}}^{(2)}$ are found by a linear regression. Lastly, we rescale ${\mathbf{b}}^{(1)},{\mathbf{b}}^{(2)}$ so that the mean weight of each ${\mathbf{b}}^{(i)}$ is equal to the average weight across all the pigs: $\tfrac{1}{9}\sum_{l=1}^{9}b^{(i)}_{l}=\tfrac{1}{9*48}\sum_{i=1}^{48}\sum_{j=1}^{9}T_{ij}$. Figure 5(a) displays the two estimated growth curves ${\mathbf{b}}^{(1)}$ and ${\mathbf{b}}^{(2)}$, overlaid on the observed growth curves. Interestingly, around week five the slope of both estimated curves change, a phenomena that may warrant further study. In Figure 5(b) we see that the fitted curve $a_{i}^{(1)}{\mathbf{b}}^{(1)}+a_{i}^{(2)}{\mathbf{b}}^{(2)}$, for pig $i$, is nearly a convex combination of ${\mathbf{b}}^{(1)}$ and ${\mathbf{b}}^{(2)}$, as the scale factor $a^{(1)}_{i}+a^{(2)}_{i}$ is close to one. The relative importance of ${\mathbf{b}}^{(1)}$ as compared to ${\mathbf{b}}^{(2)}$ noticeably varies across the pigs.
@@
One issue that has not yet been discussed in this work is that ND factorizations are not unique in general, even after accounting for scaling and permutation of the rank-one factors. Various regularization penalties or additional constraints have been introduced in the NMF literature [29, Ch 4] to select a factorization that has desirable properties. In the above analysis we have implicitly selected the minimum volume [40] ND factorization by choosing rows from ${\mathbf{T}}^{(2)}$ to serve as the ${\mathbf{b}}^{(i)}$. An alternative here is take a maximum volume factorization by choosing the ${\mathbf{b}}^{(i)}$ to be the extremal rays of the cone $\mathcal{C}(\{1,\ldots,9\})\cap\text{row}({\mathbf{T}}^{(2)})$. Either factorization will yield the same reconstruction error $\|{\mathbf{T}}-{\mathbf{T}}^{(2)}\|^{2}_{F}$, but will lead to potentially different interpretations with respect to the ${\mathbf{b}}^{(i)}$ curves and the ${\mathbf{a}}^{(i)}$ membership weights.
@@ 11.2 Perceived Mental Health of Canadians
We examine a subset of the data from the annual Canadian Community Health Survey (CCHS) [52] from 20162016-$2022$ that records responses of whether an individual perceives their mental health to be either fair or poor. The survey also records the sex and age group of the participant. The $(5,7,2)$-dimensional tensor $\mathbf{T}$} in Table 2 displays the proportion of respondents who indicated that they had either poor or fair mental health.
@@
Using the ND HALs algorithm from Section 10 we fit both rank-two and rank-one approximations to the data tensor of the form $\sum_{i=1}^{r}{\mathbf{a}}^{(i)}\otimes{\mathbf{b}}^{(i)}\otimes{\mathbf{c}}^{(i)}$. The residual sum of squares for the respective approximations are $145$ and $55$ with the total sum of squares equaling $6925$. We only impose nonnegativity constraints on the sex mode, while the Hasse diagrams in Figure 6 provides orderings on the age group and years that are consistent with the observed data. Specifically, the mental health of respondents deteriorated in 2020 and the subsequent two years, conceivably as a result of the COVID-19 pandemic. Respondents in the second youngest age group were most likely to perceive their mental health as fair or poor while respondents in the oldest age group were most likely to not provide this response.
There is a marked difference between how individuals in the 12-17 age group responded between females and males, with females having poorer perceived mental health. In fitting a rank-two ND decomposition the two rank-one factors provide a low-dimensional summary of this sex-age interaction. The rank-one factors of the two rank-one terms are shown in Table 3. These factors are scaled so that $\|{\mathbf{c}}^{(i)}\|_{1}=1$ and $\|{\mathbf{b}}^{(i)}\|_{1}=7$. We see that the factor ${\mathbf{a}}^{(2)}\otimes{\mathbf{b}}^{(2)}\otimes{\mathbf{c}}^{(2)}$ only alters the female portion of the tensor and represents a deviation between the male and female responses. In comparison with ${\mathbf{a}}^{(1)}$, the 12-17 age group in ${\mathbf{a}}^{(2)}$ takes on larger values relative to the 18-34 age group. Furthermore, the impact of pandemic is more apparent in ${\mathbf{b}}^{(2)}$ as compared to ${\mathbf{b}}^{(1)}$. In summary, the low ND rank decomposition allows for a concise interpretation of the data tensor and illustrates that young females in 2022 were particularly prone to respond that they had fair or poor mental health as compared to young males.
@@
Age group ${\mathbf{a}}^{(1)}$ ${\mathbf{a}}^{(2)}$
@@ 12 Conclusion and Future Directions
In this article the concept of the nondecreasing rank was introduced and properties of nondecreasing factorizations were examined. There remains a wealth of relevant, open questions to be explored, a few of which we mention here. One property of ND factorizations not extensively discussed is the uniqueness of such factorizations. Like with nonnegative matrix factorizations, it is expected that only under similarly strong conditions, such as separability [29, Sec 4.2.2], will a matrix have a unique ND factorization. Also not discussed, is how to choose a suitable rank $r$ for performing an ND factorization. An easy method for doing this is to plot the value of a divergence objective function against $r$ and look for a kink in the plot where the divergence stops decreasing rapidly in response to an increase of $r$ [15]. In Section 10 an algorithm is introduced for minimizing the Frobenius norm between a tensor and its ND approximation. Developing efficient multiplicative update [35] algorithms for finding ND approximations would be a valuable contribution. To construct such an algorithm more constraints as compared to NMF, stemming from the monotonicity constraints of the order cones, have to be contended with.
Nondecreasing factorizations are closely related to ideas in the field of order constrained statistical inference [7, 47]. Beyond monotonicity constraints, related shape constraints for functions, such as convexity and unimodality, would be interesting to inspect for low-rank structures. Unimodal matrix factorizations have previously been examined in [4, 13]. Unimodality is related to the umbrella order where $x_{1}\prec\cdots\prec x_{l-1}\prec x_{l}\succ x_{l+1}\cdots\succ x_{p}$. The main distinction is that in a unimodal vector it is not known which of the entries is the maximum, while for the umbrella order above the $x_{l}$th entry is posited to be the maximum. The non-convex cone of unimodal vectors is therefore equal to a union of umbrella-order cones. Problems involving convex, unimodal, and monotone factorizations can also be contextualized from an infinite-dimensional, functional data analysis perspective, where notions of convex sets and rank continue to be well-defined.
Acknowledgments